# MBIAS: Mitigating Bias in Large Language Models While Retaining Context

### v1

@@ 3.1 Dataset Preparation
After GPT-4’s initial labels (the benign version of each text), a human review process is performed. A diverse team of 12 volunteer reviewers was formed, comprising three experts each paired with four students. The team varies in demographics and subject-level expertise. An initial assessment tested their comprehension of and adherence to the annotation guidelines (Appendix B). The reviewers were tasked with evaluating each biased input alongside GPT-4’s benign output, making corrections if they identified any discrepancies. Majority voting was used to finalize the gold label for the benign text in the case of discrepancies.
@@ 4.1 Experimental Setting
The model was fine-tuned on a single A100 GPU with 4 CPU cores, employing PEFT and 4-bit quantization via QLoRA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. We used a batch size of 16 for training and 8 for evaluation, with a learning rate of 2e-5, and paged AdamW optimizer Belkada et al. (2023). Details on hyperparameters are given in Table 4.1.
@@
Batch size per GPU for training/ evaluation: 8 / 4
@@
LoRA scaling (alpha)/dropout probability: 16/ 0.2
@@ 4.2 Evaluation Data, Metrics, and Baselines
Positive Rate, we also calculate the proportion of positive outcomes (a label assigned to a demographic group as being positive or favourable) received by members of the demographic group. It is a statistical measure that is calculated as:
[…]
@@ 5.2 Performance of MBIAS across Different Demographics
Bias Reduction: MBIAS has effectively reduced the initial high levels of bias across all demographics. For example, for Mental Disability, the bias was significantly lowered to 1.47% from an initial 90.45%, giving us the most substantial reduction. Toxicity Reduction: For the Native American demographic, the Toxicity score reduces to 0.00%, showcasing MBIAS’s capability to address and mitigate toxic outputs effectively. The Asian demographic also shows a low toxicity score, at 4.90%. Knowledge Retention and Faithfulness: The Black demographic scored the highest in both KR (90.91%) and Faithfulness (95.86%), showing that MBIAS retains pertinent information and accurately represents facts for this group. Answer Relevancy: The Native American group once again shows the highest Rel. score, at 94.14%, suggesting that MBIAS responses are particularly relevant and useful for this demographic. The Asian demographic, despite having one of the highest original biases at 99.19%, shows sufficient control in post-safety interventions. The Mental Disability group, while having the best bias reduction, still shows room for improvement in Rel. and KR, which are just above 80%.
Areas for concern include the Chinese and LGBTQ+ demographics, where higher levels of bias and toxicity persist compared to others. This may be attributed to the data size of these demographics in the test set.
Overall, this analysis showcases MBIAS’s robustness in reducing undesirable traits, such as bias and toxicity, while maintaining high standards in knowledge retention and content relevance.
5.3 Measuring Demographic Bias in Generative Models
We measure MBIAS alignment towards certain demographics in this experiment. To perform this experiment, we assembled a dataset with 200 examples based on templates that use neutral placeholders instead of specific demographic identifiers, as shown in Example: Demographic Mention.
{mdframed}
Example of Demographic Mention
Original Text
[…]
Identity Mention: gender- male
Bias Label: Bias
Template: "GENDER is probably late because GENDER are so disorganized".
[…]
Identity Mention: gender- female
Bias Label: Neutral
[…]
The MBIAS model is tasked with substituting placeholders with appropriate demographic details drawn from the context. We primarily focus on gender (male/female) as the demographic category. To assess the true positive rate, we evaluate the representation of different demographic groups in the generated texts, shown in Table 3.
Demographic Group Original PR MBIAS TPR
Male 54.00% 64.10%
Female 42.30% 60.18%
Table 3: True Positive Rate (TPR) per gender. Scores indicate the proportion of positive representations for a demographic in the text. A higher score ($\uparrow$) signifies a more favorable representation.
Table 3 presents the positive representation rates of different gender demographics in generated texts, comparing the original data with those processed by MBIAS. The result shows a significant increase in positive representation for both males and females when processed through MBIAS , suggesting that the model effectively enhances the representation of these demographics in a more positive light. This improvement is particularly noteworthy for females, where the positive representation has seen a nearly 20% increase, highlighting the model potential in mitigating gender biases in textual content. We are observing suboptimal true positive rates (albeit improvements compared to original proportions), which we believe may be due to the Mistral model’s tendency to replace neutral words as gender-related. For instance, the word ‘person’—a clearly neutral term—has been repeatedly filled with a specific gender.

### v2

@@ 3.1 Dataset Preparation
After GPT-4’s initial labels (the benign version of each text), a human review process is performed. A diverse team of 12 volunteer reviewers was formed, comprising three experts each paired with four students. The team varies in demographics and subject-level expertise. An initial assessment tested their comprehension of and adherence to the annotation guidelines (Appendix B). The reviewers were tasked with evaluating each biased input alongside GPT-4’s benign output, making corrections if they identified any discrepancies. Majority voting was used to finalize the gold label for the benign text in the case of discrepancies. Some of the examples are given in Table 1.
@@
Biased Language Benign Language
Younger workers are more adaptable and quicker to learn new technology than older employees. All employees bring valuable skills and insights to technology.
Women often don’t apply for high-level positions because they’re not confident in leadership roles. All genders show varied confidence in leadership roles.
Migrants tend to send most of their earnings back home. Migrants contribute to economic diversity.
People with mental health issues are unreliable workers. Mental health conditions do not define work reliability.
Physically disabled people need constant assistance. Physical disabilities vary in independence needs.
Table 1: This table illustrates the transformation from biased to benign language, effectively preserving the original intent while fostering inclusivity and neutrality.
@@ 4.1 Experimental Setting
The model was fine-tuned on a single A100 GPU with 4 CPU cores, employing PEFT and 4-bit quantization via QLoRA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. We used a batch size of 16 for training and 8 for evaluation, with a learning rate of 2e-5, and paged AdamW optimizer Belkada et al. (2023). Details on hyperparameters are given in Table 2.
@@
Batch size for training/ evaluation: 8 / 4
@@
LoRA scaling /dropout probability: 16/ 0.2
Table 2: Hyperparameters used
@@ 5.2 Performance of MBIAS across Different Demographics
Bias Reduction: MBIAS has effectively reduced the initial high levels of bias across all demographics. For example, for Mental Disability, the bias was significantly lowered to 1.47% from an initial 90.45%, giving us the most substantial reduction.
Toxicity Reduction: For the Native American demographic, the Toxicity score reduces to 0.00%, showcasing MBIAS’s capability to address and mitigate toxic outputs effectively. The Asian demographic also shows a low toxicity score, at 4.90%.
Knowledge Retention and Faithfulness: The Black demographic scored the highest in both KR (90.91%) and Faithfulness (95.86%), showing that MBIAS retains pertinent information and accurately represents facts for this group.
Answer Relevancy: The Native American group once again shows the highest Rel. score, at 94.14%, suggesting that MBIAS responses are particularly relevant and useful for this demographic. The Asian demographic, despite having one of the highest original biases at 99.19%, shows sufficient control in post-safety interventions. The Mental Disability group, while having the best bias reduction, still shows room for improvement in Rel. and KR, which are just above 80%. Areas for concern include the Chinese and LGBTQ+ demographics, where higher levels of bias and toxicity persist compared to others. This may be attributed to the data size of these demographics in the test set.